Write the slope-intercept equation for the line that passes through
(6, -9) and has slope -7/6.
I need a little assistance, please! Also, if you could show your work, I would appreciate it. Basically, I would like to see how you arrived to your answer. Thank you very much for your time!

Answer by sudhanshu_kmr(1152)


equation..
(y - y1) = m(x-x1)
=> (y+9) = -7/6 (x-6)
=> y = -7x/6 + 7 -9
=> y = -7x/6 -2
slope = -7/6 , y-intercept = -2

if any doubt, feel free to contact..: sudhanshu.cochin@gmail.com